# Beyond Numeric Awards: In-Context Dueling Bandits with LLM Agents

### v1

@@ Appendix A Algorithm Design and Analysis of IF-Enhanced LLM
In this section, we detail the design principles and implementation of the IF-Enhanced LLM algorithm. We also provide a rigorous proof of Theorem 5.1 and 5.2, establishing the theoretical guarantees of IF-Enhanced LLM under the assumptions outlined in Appendix A.2.1.
A.1 Detailed Procedure Description
@@
In Procedure 1 below, we describe the Match Arms procedure used in IF-Enhanced LLM (see Algorithm 1 and Figure 3).
Algorithm 1 Match Arms
@@
Input: Two arms $a,a^{\prime}$, confidence parameter $\delta\leftarrow 1/(K^{2}\log T)$, and threshold $\epsilon$
@@
else return $B\leftarrow B$, $a$
return $B\leftarrow B\setminus\{b^{\prime}\}$, $b$
We reprise the IF2 procedure in [3] below to complement the presentation of IF-Enhanced LLM.
@@
Input: Incumbent arm $a$, candidate arms $B$, confidence parameter $\delta\leftarrow 1/(TK^{2})$
if $t\leq(8K/\epsilon_{1,2}^{2})\log(TK)$ then
             for $b\in B$ do  Compare $a$ with $b$ and update $\hat{P}_{a,b}$ and $\hat{C}_{a,b}$, $t\leftarrow t+1$
@@
else return $B\leftarrow B$
$B\leftarrow B\backslash\{b^{\prime}\}$ if $\exists$ $b^{\prime}\in B$ such that $\hat{P}_{a,b^{\prime}}<1/2$ and $1/2\notin\hat{C}_{a,b^{\prime}}$ then
       while $\exists b\in B$ such that $\hat{P}_{a,b}>1/2$ do  $B\leftarrow B\backslash\{b\}$
@@
return $B$
@@
A.2 Theoretical Analysis
A.2.1 Useful Assumptions and Lemmas for Dueling Bandits
We introduce the useful assumptions and lemmas for Dueling Bandits that are necessary for the theoretical analysis of our proposed algorithm.
@@ A.2.1 Useful Assumptions and Lemmas for Dueling Bandits
With the TO property satisfied, we assume the preference matrix $P$ further satisfies the following two standard properties [6, 17, 3].
@@
Note that the Bradley-Terry-Luce (BTL) model [34] used in our experiments 4.1 satisfies Assumption 3 and 4. We restate the following theoretical guarantees for IF2 that is useful in the proof of Theorem 5.2. Let $\epsilon_{\mathrm{bad}}\coloneq\min_{b\neq b^{*}}\epsilon(b,b^{*})$.
Lemma 1 (Theorem 2 in [3]).
@@
The following expected regret bound achieved by IF2 is tight up to multiplicative constants, as indicated by the lower bound (Theorem 4) in [3] such that any algorithm Alg for DB satisfies $\mathbb{E}[\mathsf{SR}(\textsc{Alg})]=\Omega\left((K/\epsilon_{\mathrm{bad}})\log T\right)$.
A.2.2 Theoretical Guarantees of IF-Enhanced LLM
@@
Consider the following Dueling Bandit instance with $K\geq 3$ arms $\{b_{1},\ldots,b_{K}\}$ and preference matrix $P$:
@@
According to the Assumption 1, in the worst case, the LLM agent’s behavior is equivalent to randomly selecting a duel in each round. For $K$ arms, there are $\binom{K}{2}=\frac{K(K-1)}{2}$ possible duel combinations. Therefore, the probability of selecting a duel containing $b^{*}$ in each round is $\frac{K-1}{\binom{K}{2}}=\frac{2}{K}$, which yields $\mathbb{E}[N(T)]=T\cdot\frac{2}{K}$. The regret bound becomes:
@@
Part II: Expected Regret Bounds of IF-Enhanced LLM
Suppose at each step $t\leq T$, aligning with the design of IF2 in [3], $\hat{P}_{t}$ is estimated such that each $\smash{\hat{P}_{i,j}}$ is the fraction of number of comparisons when $b_{i}$ was the winner out of all previous $t$ comparisons. Define a confidence interval $\smash{\hat{C}_{t}\coloneq(\hat{P}_{t}-c_{t},\hat{P}_{t}+c_{t})}$ where $\smash{c_{t}\coloneq\sqrt{\log(1/\delta)/t}}$. Before proceeding to prove Theorem 5.2, we first state a useful lemma from [3] as a result of the Hoeffding’s inequality.
Lemma 2 (Generalized Lemma 1 in [3]).
With a confidence parameter $\delta=1/(K^{2}\log T)$ with $\delta\in(0,1/2]$, a winner between two arms $b_{i}$ and $b_{j}$ is identified with probability at least $1-\delta$, using at most $\left(8/\epsilon_{i,j}^{2}\right)\log(K\log T)$ number of comparisons.
Note that Lemma 2 can be directly implied by Lemma 1 in [3]. Now, under Assumption 3 and 4 such that the preference matrix $P$ satisfies the SST and STI properties, we prove Theorem 5.2.
@@
First Bound. The first $T_{\mathrm{LLM}}$ steps induce accumulated weak/strong regret of at most $O(T_{\mathrm{LLM}})$. Applying Lemma 2, with $(8/\epsilon_{1,2}^{2})\log(K\log T)$ (by setting a hyper-parameter $\epsilon=\epsilon_{1,2}$) number of comparisons between two arms, since the best arm $b^{*}$ is always included in each comparison, the best arm $b^{*}$ is correctly identified with probability at least $1-1/(K^{2}\log T)$. This procedure leads to no weak regret since $b^{*}$ suggested by the LLM agent is always included as the incumbent arm in future comparisons. Regarding the strong regret, the union bound implies that with probability $1-1/(K\log T)$, $b^{*}$ will win all the $K+1$ matches, thus the expected strong regret induced by the steps after time $T_{\mathrm{LLM}}$ can be bounded by
@@
$\leq$ $\left(1-\frac{1}{K\log T}\right)O\left(\frac{K\log(K\log T)}{\epsilon_{1,2}}\right)+\frac{1}{K\log T}O\left(\frac{K}{\epsilon_{1,2}}{\log T}\right)$
@@
Second Bound. Furthermore, consider the adversarial selection of arms from the LLM agent. According to Lemma 2, the IF2 procedure with $\smash{O((K/\epsilon_{1,2}^{2})\log(TK))}$ is implemented at most $K$ times with probability $1-1/(TK)$, provided with $|B|=K$. Finally, each implementation of Procedure 2 costs regret (both strong and weak) of at most $O(8/\epsilon_{1,2})\log(K\log T)$ since there are at most $(8/\epsilon_{1,2})\log(K\log T$ comparisons of pairs. Applying the expected regret bound in Lemma 1 completes the proof.
Appendix B Prompt Design and Supplementary Results
B.1 LLM Experiments
@@
In this section, we provide the detailed design of the prompts used in our experiments and provide additional results to support our findings. We begin by presenting the original prompt used in the LLM-Env interaction. We then introduce the perturbed prompts, which include both noisy and adversarial variations to test the robustness of our approach. Finally, we provide a five consecutive time steps example of GPT-4 Turbo using original prompt to showcase its behavior in the dueling bandits setting.
B.1.1 Design of Prompts

### v3

@@ Appendix B Algorithm Design and Analysis of LEAD
In this section, we detail the design principles and implementation of the LEAD algorithm. First, we present the algorithm design logic. Then, we provide a rigorous proof of Theorem 4.14.2, and 4.3, establishing the theoretical guarantees of LEAD (IF2 base) under the assumptions outlined in Appendix B.3.1.
B.1 Algorithm Design Logic
Explore-then-Exploit algorithms as Ideal Candidates. Classic DB algorithms can be classified into two categories: Explore-Then-Exploit methods and Ongoing Regret Minimization methods [Zoghi et al., 2014a]. Among these, Explore-Then-Exploit structure stands out as particularly well-suited for LLM augmentation:
  • The Explore-Then-Exploit structure naturally aligns with the LLMs’ tendency to keep exploring without converging (see Figure 3), allowing for leveraging the LLMs’ exploration behavior while mitigating their exploration vulnerability and convergence instability (see Section 3.2).
  • Its symbolic representation of the algorithm’s logic enables clear integration of LLM suggestions at specific points without disrupting the overall structure and theoretical guarantees. In contrast, algorithms like Self-Sparring in [Sui et al., 2017] are less symbolic, making them less suitable for direct LLM augmentation.
  • Its strong theoretical guarantees, e.g., IF2 with an expected regret bound of $O((K/\epsilon_{\mathrm{bad}})\log T)$ matching the DB problem’s lower bound of $\Omega((K/\epsilon_{\mathrm{bad}})\log T)$ up to constants (see Appendix B.3.1), and its empirical performance (see Figures 2 and 9) provide a robust foundation, ensuring convergence and bounded regret.
B.2 Detailed Procedure Description
@@
In Procedure 1 below, we describe the Match Arms procedure used in LEAD (see Algorithm 1 and Figure 4).
Algorithm 1 Match Arms (with a bounded number of comparisons)
@@
Input: Two arms $a,a^{\prime}$, confidence parameter $\delta\leftarrow 1/(K^{2}\log T)$, and threshold $\epsilon\leftarrow\epsilon_{1,2}$
@@
      return $b$
else return $a$
Algorithm 2 Validate
Input: Incumbent arm $a$, candidate arms $B$, $\mathsf{TrustLLM}$, confidence parameter $\delta\leftarrow 1/(TK^{2})$, and threshold $\epsilon\leftarrow\epsilon_{1,2}$
if $\mathsf{TrustLLM}$ is $\mathtt{True}$ then
       for $b\in B$ do
             if $t\leq(16/\epsilon^{2})\log(K\log T)$ then
                         Compare $a$ with $b$ and update $\hat{P}_{a,b}$ and $\hat{C}_{a,b}$, $t\leftarrow t+1$
                   end while
             if $b\neq a$ then return $\mathsf{StillTrust}\leftarrow\mathtt{False},B\leftarrow B\backslash\{a\}$
       end for
      return $\mathsf{StillTrust}\leftarrow\mathtt{True},B\leftarrow\emptyset$
if $\mathsf{TrustLLM}$ is $\mathtt{False}$ then return $\mathsf{StillTrust}\leftarrow\mathtt{False},B\leftarrow B$
We reprise the IF2 procedure in [Yue et al., 2012] below to complement the presentation of LEAD.
@@
Input: Incumbent arm $a$, candidate arms $B$, confidence parameter $\delta\leftarrow 1/(TK^{2})$, $t\leftarrow 0$
if $t\leq(16K/\epsilon_{1,2}^{2})\log(K\log T)$ then
       for $b\in B$ do
             Compare $a$ with $b$ and update $\hat{P}_{a,b}$ and $\hat{C}_{a,b}$, $t\leftarrow t+1$
       end for
      $a,B\leftarrow\textsc{Anneal}(a,B)$
return $a,B$
Algorithm 4 Anneal
Input: Incumbent arm $a$, candidate arms $B$, confidence parameter $\delta\leftarrow 1/(TK^{2})$, matrices $\hat{P}$ and $\hat{C}$
while $\exists\ (b,b^{\prime})\in B$ such that $\hat{P}_{b,b^{\prime}}>1/2$ and $1/2\notin\hat{C}_{b,b^{\prime}}$ do
       $B\leftarrow B\backslash\{b^{\prime}\}$
@@
if $\exists$ $b^{\prime}\in B$ such that $\hat{P}_{a,b^{\prime}}<1/2$ and $1/2\notin\hat{C}_{a,b^{\prime}}$ then
       while $\exists b\in B$ such that $\hat{P}_{a,b}>1/2$ do
             $B\leftarrow B\backslash\{b\}$
@@
       end while
      $a\leftarrow b^{\prime}$, $B\leftarrow B\backslash\{b^{\prime}\}$
return $a,B$
@@
B.3 Theoretical Analysis
B.3.1 Useful Assumptions and Lemmas for Dueling Bandits
We introduce the useful definitions, assumptions and lemmas for Dueling Bandits that are necessary for the theoretical analysis of our proposed algorithm.
Throughout this paper, we consider two important performance metrics. The first is the strong regret of a given algorithm ALG, defined as
which only compares $b^{*}$ against the better of the two selected arms $\mathsf{Arm}_{1}(t)$ and $\mathsf{Arm}_{2}(t)$. It is worth highlighting that LLM agents exhibit significantly different behaviors with respect to the two defined notions of regret, as detailed in Section 3.2.
@@ B.3.1 Useful Assumptions and Lemmas for Dueling Bandits
With the TO property satisfied, we assume the preference matrix $P$ further satisfies the following two standard properties [Yue and Joachims, 2009, 2011, Yue et al., 2012].
@@
Note that the Bradley-Terry-Luce (BTL) model [Bradley and Terry, 1952] used in our experiments 3.1 satisfies Assumption 2 and 3. We restate the following theoretical guarantees for IF2 that is useful in the proof of Theorem 4.2. Let $\epsilon_{\mathrm{bad}}\coloneq\min_{b\neq b^{*}}\epsilon(b,b^{*})$.
Lemma 1 (Theorem 2 in [Yue et al., 2012]).
@@
The following expected regret bound achieved by IF2 is tight up to multiplicative constants, as indicated by the lower bound (Theorem 4) in [Yue et al., 2012] such that any algorithm Alg for DB satisfies $\mathbb{E}[\mathsf{SR}(\textsc{Alg})]=\Omega\left((K/\epsilon_{\mathrm{bad}})\log T\right)$.
B.3.2 Theoretical Guarantees of LEAD
@@
Assumption 4 (Worst-Case Behavior).
Under the original prompt (see Figure 7), the worst-case behavior of an LLM agent in the dueling bandit setting is equivalent to a randomizer that selects action pairs uniformly at random.
Vulnerability of Standalone LLM Agents. Inspired by the adversarial corruptions framework introduced in [Hajiesmaili et al., 2020] for the classic MAB problem, we investigate the vulnerability of standalone LLM agents in the DB setting under adversarial prompts. We consider an attacker with a budget $\Phi(T)$ who employs the following strategy: whenever the LLM agent selects the optimal arm $b^{*}$ for comparison, the attacker manipulates the input prompt to the LLM to eliminate $b^{*}$ from the duel with probability $p$ (where $0<p\leq 1$ is a constant), subject to the constraint of performing at most $\Phi(T)$ attacks over $T$ rounds. This adversarial strategy compels the LLM agent to select suboptimal arms, resulting in poor performance, as formalized in the following theorem with Assumption 4.
@@
Consider the following DB instance with $K\geq 3$ arms $\{b_{1},\ldots,b_{K}\}$ and preference matrix $P$:
@@
According to the Assumption 4, in the worst case, the LLM agent’s behavior is equivalent to randomly selecting a duel in each round. For $K$ arms, there are ${K(K-1)}/{2}$ possible duel combinations. Therefore, the probability of selecting a duel containing $b^{*}$ in each round is $(K-1)/{\binom{K}{2}}=\frac{2}{K}$, which yields $\mathbb{E}[N(T)]=T\cdot\frac{2}{K}$. The regret bound becomes:
@@
Part II: Expected Regret Bounds of LEAD (IF2 base)
Suppose at each step $t\leq T$, aligning with the design of IF2 in [Yue et al., 2012], $\hat{P}_{t}$ is estimated such that each $\smash{\hat{P}_{i,j}}$ is the fraction of number of comparisons when $b_{i}$ was the winner out of all previous $t$ comparisons. Define a confidence interval $\smash{\hat{C}_{t}\coloneq(\hat{P}_{t}-c_{t},\hat{P}_{t}+c_{t})}$ where $\smash{c_{t}\coloneq\sqrt{\log(1/\delta)/t}}$. Before proceeding to prove Theorem 4.2, we first state a useful lemma from [Yue et al., 2012] as a result of the Hoeffding’s inequality [Hoeffding, 1994].
Lemma 2 (Generalized Lemma 1 in [Yue et al., 2012]).
Let $\delta=1/(K\log T)^{2}$ be a confidence parameter with $\delta\in(0,1/2]$, a winner between two arms $b_{i}$ and $b_{j}$ is identified with probability at least $1-\delta$, using at most $\left(16/\epsilon_{i,j}^{2}\right)\log(K\log T)$ number of comparisons.
Note that Lemma 2 can be directly implied by Lemma 1 in [Yue et al., 2012]. Now, under Assumption 2 and 3 such that the preference matrix $P$ satisfies the SST and STI properties, we prove Theorem 4.2.
@@
Weak Regret Bound. The first $T_{\mathrm{LLM}}$ steps induce accumulated weak regret of at most $O(T_{\mathrm{LLM}})$. According to [Yue et al., 2012], IF2 plays $O(K)$ matches (comparisons) in expectation. Thus, the expected number of rounds of calling IF2 Procedure is $O(\log T/\log(K\log T))$. Applying Lemma 2, with $\smash{O\left((1/\epsilon_{1,2}^{2})\log(K\log T)\right)}$ (by setting a hyper-parameter $\epsilon=\epsilon_{1,2}$) comparisons between two arms, since the best arm $b^{*}$ is always included in each comparison, the best arm $b^{*}$ is correctly identified with probability at least $1-1/(K\log T)^{2}$. This procedure leads to no weak regret since $b^{*}$ suggested by the LLM agent is always included as the incumbent arm in future comparisons.
Moreover, the implementation of Procedure 3 induces at most $O((K/\epsilon_{1,2}^{2})\log(K\log T))$ comparisons. The validation procedure (Procedure 2) leads to no weak regret if $b_{\mathrm{LLM}}$ is indeed the best arm and the identification of $b_{\mathrm{LLM}}=b^{*}$ succeeds with a probability $1-1/T$. Denote by $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ two error events when $b^{*}$ loses some of the matches in the LLM Phase. there exist comparisons (matches) fail in the validation procedure (Procedure 2) or the IF2 Phase (Procedure 3). The union bound implies with probability $1-1/(K\log T)$, $b^{*}$ will win all the matches such that $P(\mathcal{E}_{1})\leq 1/(K\log T)$. Similarly, $P(\mathcal{E}_{2})\leq 1/T$. Combining these events, regarding the total expected weak regret, the expected weak regret induced by the steps after time $T_{\mathrm{LLM}}$ can be bounded by
@@
[…]
@@
Convergence Guarantee. Furthermore, consider the adversarial selection of arms from the LLM agent. According to Lemma 2, the IF2 procedure with an expected regret $\smash{O\left((K/\epsilon_{1,2})\log(T)\right)}$ is implemented at most $O(1)$ times with probability $1-1/(TK)$, provided with $|B|=K$. Therefore, the expected regret (either strong or weak) induced by each implementation of Procedure 3 is at most $O\left((K/\epsilon_{1,2})\log(T)\right)$ since there are at most $O\left((K/\epsilon_{1,2}^{2})\log(K\log T)\right)$ additional comparisons of pairs in the LLM phase. Finally, applying the expected regret bound in Lemma 1 completes the proof.
Part III: Converse
In the following, we argue that for any algorithm ALG, achieving an upper bound $\mathbb{E}\left[\mathsf{WR}(\textsc{ALG})\right]\leq T_{\mathrm{LLM}}$ for all $T_{\mathrm{LLM}}$ is impossible.
Proof of Theorem 4.3.
Suppose ALG is an algorithm that leads to a weak regret bound $\mathbb{E}\left[\mathsf{WR}(\textsc{ALG})\right]\leq T_{\mathrm{LLM}}$ for all $T_{\mathrm{LLM}}$, then it has to trust and include the recommended arm in all the comparisons immediately after it is proposed by the LLM agent to ensure that future weak regret becomes zero. To see this, note that one can always construct an adversarial $T_{\mathrm{LLM}}$ that leads to a nonzero future weak regret. However, the LLM agent can choose to provide an arm that is always not the best arm for all $t\in\{1,\ldots,T\}$. This leads to $\mathbb{E}\left[\mathsf{SR}(\textsc{ALG})\right]\geq\mathbb{E}\left[\mathsf{WR}(\textsc{ALG})\right]\geq\Omega\left(T\right)$.
Appendix C Prompt Design and Supplementary Results
C.1 LLM Experimental Results
@@
In this section, we provide the detailed design of the prompts used in our experiments and provide additional results to support our findings. We begin by presenting the original prompt used in the LLM-Env interaction and introduce the perturbed prompts, which include both noisy and adversarial variations to test the robustness of our approach. Finally, we provide four exemplars using the original prompt to to showcase the behavior of both GPT-4 Turbo and o1-preview.
C.1.1 Environments
[…]
In transitive instances, the preference matrices are constructed using the Bradley-Terry-Luce (BTL) model [Bradley and Terry, 1952, Yue et al., 2012], with a generalized form known as the Plackett-Luce model [Plackett, 1975]. In this model, each arm is associated with a utility parameter $\theta(i)>0$, where $i$ represents the rank of the arm (i.e., $\theta(1)$ corresponds to the best arm, $\theta(2)$ corresponds to the second best arm, and so on). For any pair of arms $b_{i}$ and $b_{j}$, the probability of $b_{i}$ being preferred over $b_{j}$ is determined by $P{(i\succ j)}=\theta(i)/(\theta(i)+\theta(j))$. Setting the number of arms $K=5$, we randomize the order of the arms to prevent selection bias, resulting in the following arm ordering: $b_{5}\succ b_{3}\succ b_{2}\succ b_{1}\succ b_{4}$. We use two instances: Transitive-Easy and Transitive-Hard, with their respective $\theta$ parameters given by:
  • Transitive-Easy instance: $\theta(1)=1,\ \theta(i)=0.5-(i-1)/2K$,  $\forall i\in[2,K]$.
  • Transitive-Hard instance: $\theta(i)=1-(i-1)/K$,  $\forall i\in[K]$.
Note that the datasets generated in this way satisfy the Strong Stochastic Transitivity (SST) and Stochastic Triangle Inequality (STI) properties [Yue et al., 2012] (see Appendix B.3.1 for more details). The settings of the used BTL model also imply the existence of a Condorcet winner.
Intransitive Case: $CW\setminus(\text{SST}\cup\text{STI})$
In intransitive instances, the preference matrices are constructed to violate both the Strong Stochastic Transitivity (SST) and Stochastic Triangle Inequality (STI) properties. This design creates cyclic preferences among the non-winning arms while preserving the existence of a Condorcet winner. Setting $K=5$, we still use the same shuffled arm ordering: $b_{5}\succ b_{3}\succ b_{2}\succ b_{1}\succ b_{4}$ for intransitive instances.
  • Intransitive-Easy instance: The Condorcet winner $b_{5}$ has a strong preference over any other arm:
    $P(5\succ j)=0.8,\quad P(j\succ 5)=0.2,\quad\forall j\in\{1,\dots,4\}.$
    Among the non-winning arms $b_{1},\dots,b_{4}$, cyclic preferences are introduced via:
    $P(i\succ j)=0.8-0.2\cdot\left((j-i-1)\bmod(K-1)\right),\quad\forall i,j\in\{1,\dots,4\},\ i\neq j.$
    This configuration ensures a clear dominance by $b_{5}$.
  • Intransitive-Hard instance: The Condorcet winner’s preference is weaker, with:
    $P(5\succ j)=0.6,\quad P(j\succ 5)=0.4,\quad\forall j\in\{1,\dots,4\}.$
    This setting makes it more challenging to identify $b_{5}$ as the Condorcet winner.
Finally, in both instances, the symmetry condition is imposed for consistency:
$P(j\succ i)=1-P(i\succ j),\quad\forall i,j\in\{1,\dots,K\},\ i\neq j.$
Accordingly, as shown below, we create a cyclic pattern of preferences among the non-winning arms while maintaining the Condorcet winner’s superiority.
Intransitive-Easy Instance ($p_{w}=0.8$)
$P=\begin{bmatrix}0.0&0.8&0.6&0.4&0.2\\ 0.2&0.0&0.8&0.6&0.2\\ 0.4&0.2&0.0&0.8&0.2\\ 0.6&0.4&0.2&0.0&0.2\\ 0.8&0.8&0.8&0.8&0.0\\ \end{bmatrix}$
Intransitive-Hard Instance ($p_{w}=0.6$)
$P=\begin{bmatrix}0.0&0.8&0.6&0.4&0.4\\ 0.2&0.0&0.8&0.6&0.4\\ 0.4&0.2&0.0&0.8&0.4\\ 0.6&0.4&0.2&0.0&0.4\\ 0.6&0.6&0.6&0.6&0.0\\ \end{bmatrix}$
C.1.2 Design of Prompts